# Tutorial Debriefing: Applied Statistical Causal Inference in Requirements Engineering

### v1

@@ 1 Resources
The material from our tutorial remains available under an open source license on GitHub. [^11] https://github.com/JulianFrattini/bda4sci It includes the presentation slides as well as code examples to recreate figures and analyses.
@@ 2 History
Finally, we summarize the history of the tutorial series that aims to bring SCI knowledge to the SE community. Spiritual predecessors were a similar tutorial held by Richard Torkar and Robert Feldt at ICSE’21[torkar2021bayesian] and by Robert Feldt during ESEM’22.121212https://conf.researchr.org/details/esem-2022/esem-2022-eseiw-isern/ Based on this prior work, we established the following tutorial series:
@@
Future updates will be documented in an revised version of this document.

### v2

@@ 1 Resources
The material from our tutorial remains available under an open source license on GitHub [^11] https://github.com/JulianFrattini/bda4sci and Zenodo. [^12] https://doi.org/10.5281/zenodo.17649808 It includes the presentation slides as well as code examples to recreate figures and analyses.
@@ 2 History
Finally, we summarize the history of the tutorial series that aims to bring SCI knowledge to the SE community. Spiritual predecessors were a similar tutorial held by Richard Torkar, Carlo Furia, and Robert Feldt at ICSE’21[torkar2021bayesian] and by Robert Feldt during ESEM’22.131313https://conf.researchr.org/details/esem-2022/esem-2022-eseiw-isern/ Based on this prior work, we established the following tutorial series:
@@
Future updates will be documented in an revised version of this document. The beautiful plots were created by Michael Dorner  [^15] https://www.michaeldorner.de/ using tikz.